1) Derive the expression of workdone by the position on the cylinder in terms of the pressure and the volume of fluid in the cylinder

2) A unit of a fluid at a pressure of 3 bar and with a specific volume of 0.18m3/kg contained in a cylinder behind a piston that expands reversibly to a pressure of 0.6 bar the equation of pressure is taken as below p= c/v2 where c is a constant. Calculate the workdone during the process.


1
Expert's answer
2021-10-26T10:01:14-0400

1) The work is force times change in position caused by the action of the force, right?


"W=F\Delta h.\\
F=pA,"

where A - area:


"W=pA\Delta h=p\Delta V."

2) The work is


"W=c\int^{v_2}_{v_1}\frac{dv}{v^2}=c\bigg(\frac1{v_1}-\frac1{v_2}\bigg)."

Since the product of pressure and specific volume is constant:


"c=p_1v_1^2=p_2v_2^2,\\c=3·0.18^2=0.0972\text{ bar}·(\text{ m}^3/\text{kg})^2.\\\space\\
v_2=\sqrt{c/p^2}=\sqrt{0.0972/0.6}=0.402\text{ m}^3/\text{kg}."

Therefore:

"W=0.0972·10^5\bigg(\frac1{0.18}-\frac1{0.402}\bigg)=29821\text{ J}."